# Large Language Models for Information Retrieval: A Survey

### v1

@@ 6.3 Limitations
Several IR systems applying the retrieval-augmented generation strategy, such as New Bing and Langchain, have already entered commercial use. However, there are also some challenges in this novel retrieval-augmented content generation system. These include challenges such as effective query reformulation, optimal retrieval frequency, correct document comprehension, accurate passage extraction, and effective content summarization. Addressing these challenges is crucial to fully realizing the potential of this approach in practical applications.
@@ 7 Future Direction
In this survey, we have conducted a comprehensive review of recent advancements for LLM-enhanced IR systems. Since the integration of LLMs into IR systems is still in its early stages, there are still many opportunities and challenges. In this section, we discuss some potential future directions. The future directions are organized according to the taxonomy presented in a recent perspective paper [50].
7.1 Indexing
As elaborated in Section 4.2.2, the emergence of LLMs has paved the way for generative retrievers to generate document identifiers for retrieval tasks. In this manner, the document index or knowledge about documents is encoded in the parameters of LLMs. Yet, the intrinsic static nature of LLM parameters, coupled with the expensive fine-tuning process, poses challenges for generative retrievers to update document indexes for new documents. Therefore, there is a need to explore methods for constructing an incremental index for generative retrieval.
Furthermore, the contents on the web can be multi-modal, covering diverse users’ needs, including texts, images, audios, and videos. However, previous works primarily utilized the semantic information of candidates as document identifiers and focused on text-based retrieval (uni-modal). As multi-modal LLMs continue to evolve, the design and enablement of LLMs for effective multi-modal retrieval becomes a challenge that merits thorough investigation.
7.2 User Modeling
User modeling in IR refers to the process of understanding user’s queries, preferences, and search behaviors to enhance the search experience. Due to the excellent capability of LLMs in understanding natural language, the integration of LLMs into user modeling holds significant promise for refining personalized search and interaction. Several promising directions are highlighted:
(1) Enhancing query understanding. User queries tend to be ambiguous or multi-faceted, hindering the search engine from understanding the user’s real search intent. By analyzing the query text, LLMs could identify the disambiguated terms in the query or the potential sub-intents [151], which contributes to clarifying the user’s information needs and inferring user intent more accurately.
(2) Improving personalized search. LLMs can assist in personalized searches by analyzing user-specific information. By incorporating the user’s search history (e.g., issued queries, click-through behaviors, and dwell time), LLMs can construct accurate user profiles and rerank the search results accordingly, providing better results.
(3) Advancing conversational search interfaces. LLMs can play a pivotal role in refining conversational search interfaces, providing more natural and intuitive human-machine interactions. By incorporating historical interactive information, LLMs can adapt system responses based on user preferences, providing a more personalized and effective conversational experience.
By leveraging the latent power of LLMs, these directions pave the way for more sophisticated and tailored user modeling in IR. As LLMs continue to evolve, their potential to revolutionize user-centric search paradigms remains important.
7.3 Matching or Ranking
@@
Leveraging LLMs to improve ranking models in IR has received considerable attention, promising an enhanced understanding of queries and documents for improved ranking performance. However, despite strides in this field, several limitations and challenges still need to be solved:
(1) Mitigating LLM latency. LLMs, with their massive parameters and world knowledge, often entail high latency during online applications. Given that search engines require in-time responses, balancing the efficiency of LLM-based ranking models with the timely generation of ranking lists arises as a crucial challenge.
(2) Tailoring LLMs for ranking tasks. Existing studies have demonstrated that LLMs are able to generalize to various downstream tasks in zero- or few-shot settings. However, few-shot LLMs still perform worse than the models fine-tuned on the task-specific training data [105, 107] because they lack task-specific knowledge. Consequently, it is necessary to fine-tune LLMs on ranking tasks to improve their in-domain ability. Moreover, the fine-tuning of LLMs with over 10B parameters remains under-explored due to resource constraints. Existing methods mostly concentrate on LLMs with relatively modest parameters [110], such as T5. Nevertheless, many studies confirm that larger parameters can result in exponential performance improvement [113, 119]. Consequently, the way to fine-tune (or pre-train from scratch) LLMs with tens or hundreds of billions of parameters for ranking tasks persists as an open and pressing problem.
(3) Evolution of the IR framework. Beyond employing LLMs as ranking models, a more significant problem brought by the appearance of LLMs is: Is the conventional IR framework necessary in the era of LLMs? Traditional IR aims to return a ranking list of documents that are relevant to issued queries. However, the development of generative language models has introduced a novel paradigm: the direct generation of answers to input questions by LLMs. Evidently, a single LLM is hard to ensure the timeliness, reliability, and reality of answers. Thus, it is indispensable to use IR models to provide reliable references for LLMs, resulting in two further questions: what are good references for LLMs to generate answers, and how to leverage the references provided by IR models to generate better answers.
7.4 Evaluations
@@
LLMs have attracted significant attention in the field of IR due to their strong ability in context understanding and text generation. To validate the effectiveness of LLM-based IR approaches, it is crucial to develop appropriate metrics and datasets. Traditional evaluation metrics primarily focus on comparing the retrieval results of IR models with ground truth, such as Precision, Recall, Mean Reciprocal Rank (MRR) [152], Mean Average Precision (MAP), and Normalized Discounted Cumulative Gain (nDCG) [153]. However, the widespread adoption of LLMs in IR has led to a notable enhancement in their generation capabilities. Consequently, there is an imperative demand for novel evaluation strategies to effectively assess the performance of LLM-based models in IR. Previous evaluation metrics of text generation have several limitations, including: (1) Dependency on lexical matching: Methods like BLEU [154] primarily evaluate the quality of generated outputs based on n-gram matching. This approach doesn’t account for lexical diversity and contextual semantics. As a result, models may favor generating common phrases or sentence structures rather than producing creative and novel content. (2) Insensitivity to subtle differences: Traditional evaluation methods may not be sensitive to subtle differences in generated outputs. For example, if a generated output has minor semantic differences from the reference answer but is otherwise similar, traditional methods might not capture these nuanced distinctions. How to evaluate the quality of generated content remains a challenge.
7.5 Interaction
With the increasing capabilities of LLMs, the future interaction between users and IR systems will be significantly changed. Due to the powerful natural language processing and understanding capabilities of LLMs, the traditional keyword-based search paradigm is expected to be progressively replaced by conversational search. LLMs will facilitate full-document comprehension, enabling users to ask complex, context-aware questions and receive comprehensive answers. Through the use of external plugins, LLMs have the potential to understand image and speech inputs, thereby facilitating complex multi-modal and cross-modal retrieval. Furthermore, LLMs can proactively provide information and recommendations based on users’ historical search behaviors and preferences, enhancing the overall search experience. These applications demonstrate the potential of LLMs to revolutionize the interaction paradigm of existing IR systems, thereby providing users with a more personalized and seamless experience in the future.
@@ 8 Conclusion
In this survey, we have conducted a thorough exploration of the transformative impact of Large Language Models (LLMs) on Information Retrieval (IR) across various dimensions. We have organized existing approaches into distinct categories based on their functions: query rewriting, retrieval, reranking, and reader modules. In the domain of query rewriting, LLMs have demonstrated their effectiveness in understanding ambiguous or multi-faceted queries, enhancing the accuracy of intent identification. In the context of retrieval, LLMs have improved retrieval accuracy by enabling more nuanced matching between queries and documents, considering context as well. Within the reranking realm, LLM-enhanced models consider more fine-grained linguistic nuances when re-ordering results. The incorporation of reader modules in IR systems represents a significant step towards generating comprehensive responses instead of mere document lists. The integration of LLMs into IR systems has brought about a fundamental change in how users engage with information and knowledge. From query rewriting to retrieval, reranking, and reader modules, LLMs have enriched each aspect of the IR process with advanced linguistic comprehension, semantic representation, and context-sensitive handling. As this field continues to progress, the journey of LLMs in IR portends a future characterized by more personalized, precise, and user-centric search encounters.

### v2

@@ 6.3 Limitations
Several IR systems applying the retrieval-augmented generation strategy, such as New Bing and Langchain, have already entered commercial use. However, there are also some challenges in this novel retrieval-augmented content generation system. These include challenges such as effective query reformulation, optimal retrieval frequency, correct document comprehension, accurate passage extraction, and effective content summarization. It is crucial to address these challenges in order to effectively realize the potential of LLMs in this paradigm.
@@ 7 Future Direction
In this survey, we comprehensively reviewed recent advancements in LLM-enhanced IR systems and discussed their limitations. Since the integration of LLMs into IR systems is still in its early stages, there are still many opportunities and challenges. In this section, we summarize the potential future directions in terms of the four modules in an IR system we just discussed, namely query rewriter, retriever, reranker, and reader. In addition, as evaluation has also emerged as an important aspect, we will also introduce the corresponding research problems that need to be addressed in the future. Another discussion about important research topics on applying LLMs to IR can be found in a recent perspective paper [53].
7.1 Query Rewriter
LLMs have enhanced query rewriting for both ad-hoc and conversational search scenarios. Most of the existing methods rely on prompting LLMs to generate new queries. While yielding remarkable results, the refinement of rewriting quality and the exploration of potential application scenarios require further investigation.
$\bullet$ Rewriting query according to ranking performance. A typical paradigm of prompting-based methods is providing LLMs with several ground-truth rewriting cases (optional) and the task description of query rewriting. Despite LLMs being capable of identifying potential user intents of the query [153], they lack awareness of the resulting retrieval quality of the rewritten query. The absence of this connection can result in rewritten queries that seem correct yet produce unsatisfactory ranking results. Although some existing studies have used reinforcement learning to adjust the query rewriting process according to generation results [94], a substantial realm of research remains unexplored concerning the integration of ranking results.
$\bullet$ Improving query rewriting in conversational search. As yet, primary efforts have been made to improve query rewriting in ad-hoc search. In contrast, conversational search presents a more developed landscape with a broader scope for LLMs to contribute to query understanding. By incorporating historical interactive information, LLMs can adapt system responses based on user preferences, providing a more effective conversational experience. However, this potential has not been explored in depth. In addition, LLMs could also be used to simulate user behavior in conversational search scenarios, providing more training data, which are urgently needed in current research.
$\bullet$ Achieving personalized query rewriting. LLMs offer valuable contributions to personalized search through their capacity in analyzing user-specific data. In terms of query rewriting, with the excellent language comprehension ability of LLMs, it is possible to leverage them to build user profiles based on users’ search histories (e.g., issued queries, click-through behaviors, and dwell time). This empowers the achievement of personalized query rewriting for enhanced information retrieval and finally benefits personalized search or personalized recommendation.
7.2 Retriever
@@
Leveraging LLMs to improve retrieval models has received considerable attention, promising an enhanced understanding of queries and documents for improved ranking performance. However, despite strides in this field, several challenges and limitations still need to be investigated in the future:
$\bullet$ Reducing the latency of LLM-based retrievers. LLMs, with their massive parameters and world knowledge, often entail high latency during the inferring process. This delay poses a significant challenge for practical applications of LLM-based retrievers, as search engines require in-time responses. To address this issue, promising research directions include transferring the capabilities of LLMs to smaller models, exploring quantization techniques for LLMs in IR tasks, and so on.
$\bullet$ Simulating realistic queries for data augmentation. Since the high latency of LLMs usually blocks their online application for retrieval tasks, many existing studies have leveraged LLMs to augment training data, which is insensitive to inference latency. Existing methods that leverage LLMs for data augmentation often generate queries without aligning them with real user queries, leading to noise in the training data and limiting the effectiveness of retrievers. As a consequence, exploring techniques such as reinforcement learning to enable LLMs to simulate the way that real queries are issued holds the potential for improving retrieval tasks.
$\bullet$ Incremental indexing for generative retrieval. As elaborated in Section 4.2.2, the emergence of LLMs has paved the way for generative retrievers to generate document identifiers for retrieval tasks. This approach encodes document indexes and knowledge into the LLM parameters. However, the static nature of LLM parameters, coupled with the expensive fine-tuning costs, poses challenges for updating document indexes in generative retrievers when new documents are added. Therefore, it is crucial to explore methods for constructing an incremental index that allows for efficient updates in LLM-based generative retrievers.
$\bullet$ Supporting multi-modal search. Web pages usually contain multi-modal information, including texts, images, audios, and videos. However, existing LLM-enhanced IR systems mainly support retrieval for text-based content. A straightforward solution is to replace the backbone with multi-modal large models, such as GPT-4 [80]. However, this undoubtedly increases the cost of deployment. A promising yet challenging direction is to combine the language understanding capability of LLMs with existing multi-modal retrieval models. By this means, LLMs can contribute their language skills in handling different types of content.
7.3 Reranker
In section 5, we have discussed the recent advanced techniques of utilizing LLMs for the reranking task. Some potential future directions in reranking are discussed as follows.
$\bullet$ Enhancing the online availability of LLMs. Though effective, many LLMs have a massive number of parameters, making it challenging to deploy them in online applications. Besides, many reranking methods [130, 131] rely on calling LLM APIs, incurring considerable costs. Consequently, devising effective approaches (such as distilling to small models) to enhance the online applicability of LLMs emerges as a research direction worth exploration.
$\bullet$ Improving personalized search. Many existing LLM-based reranking methods mainly focus on the ad-hoc reranking task. However, by incorporating user-specific information, LLMs can also improve the effectiveness of the personalized reranking task. For example, by analyzing users’ search history, LLMs can construct accurate user profiles and rerank the search results accordingly, providing personalized results with higher user satisfaction.
$\bullet$ Adapting to diverse ranking tasks. In addition to document reranking, there are also other ranking tasks, such as response ranking, evidence ranking, entity ranking and etc., which also belong to the universal information access system. Navigating LLMs towards adeptness in these diverse ranking tasks can be achieved through specialized methodologies, such as instruction tuning. Exploring this avenue holds promise as an intriguing and valuable research trajectory.
7.4 Reader
With the increasing capabilities of LLMs, the future interaction between users and IR systems will be significantly changed. Due to the powerful natural language processing and understanding capabilities of LLMs, the traditional search paradigm of providing ranking results is expected to be progressively replaced by synthesizing conclusive answering passages for user queries using the reader module. Although such strategies have already been investigated by academia and facilitated by industry as we stated in Section V, there still exists much room for exploration.
$\bullet$ Improving the reference quality for LLMs. To support answer generation, existing approaches usually directly feed the retrieved documents to the LLMs as references. However, since a document usually covers many topics, some passages in it may be irrelevant to the user queries and can introduce noise during LLMs’ generation. Therefore, it is necessary to explore techniques for extracting relevant snippets from retrieved documents, enhancing the performance of retrieval-augmented generation.
$\bullet$ Improving the answer reliability of LLMs. Incorporating the retrieved references has significantly alleviated the hallucination problem of LLMs. However, it remains uncertain whether the LLMs refer to these supported materials during answering queries. Some studies [154] have revealed that LLMs can still provide unfaithful answers even with additional references. Therefore, the reliability of the conclusive answers might be lower compared to the ranking results provided by traditional IR systems. It is essential to investigate the influence of these references on the generation process, thereby improving the credibility of reader-based novel IR systems.
7.5 Evaluation
LLMs have attracted significant attention in the field of IR due to their strong ability in context understanding and text generation. To validate the effectiveness of LLM-enhanced IR approaches, it is crucial to develop appropriate evaluation metrics. Given the growing significance of readers as integral components of IR systems, the evaluation should consider two aspects: assessing ranking performance and evaluating generation performance.
$\bullet$ Generation-oriented ranking evaluation. Traditional evaluation metrics for ranking primarily focus on comparing the retrieval results of IR models with ground-truth (relevance) labels. Typical metrics include precision, recall, mean reciprocal rank (MRR) [155], mean average precision (MAP), and normalized discounted cumulative gain (nDCG) [156]. These metrics measure the alignment between ranking results and human preference on using these results. Nevertheless, these metrics may fall short in capturing a document’s role in the generation of passages or answers, as their relevance to the query alone might not adequately reflect this aspect. This effect could be leveraged as a means to evaluate the usefulness of documents more comprehensively. A formal and rigorous evaluation metric for ranking that centers on generation quality has yet to be defined.
@@ 7.5 Evaluation
$\bullet$ Text generation evaluation. The wide application of LLMs in IR has led to a notable enhancement in their generation capability. Consequently, there is an imperative demand for novel evaluation strategies to effectively evaluate the performance of passage or answer generation. Previous evaluation metrics for text generation have several limitations, including: (1) Dependency on lexical matching: methods such as BLEU [157] or ROUGE [158] primarily evaluate the quality of generated outputs based on $n$-gram matching. This approach cannot account for lexical diversity and contextual semantics. As a result, models may favor generating common phrases or sentence structures rather than producing creative and novel content. (2) Insensitivity to subtle differences: existing evaluation methods may be insensitive to subtle differences in generated outputs. For example, if a generated output has minor semantic differences from the reference answer but is otherwise similar, traditional methods might overlook these nuanced distinctions. (3) Lack of ability to evaluate factuality: LLMs are prone to generating “hallucination” problems [159, 160, 161]. The hallucinated texts can closely resemble the oracle texts in terms of vocabulary usage, sentence structures, and patterns, while having non-factual content. Existing methods are hard to identify such problems, while the incorporation of additional knowledge sources such as knowledge bases or reference texts could potentially aid in addressing this challenge.
@@ 8 Conclusion
In this survey, we have conducted a thorough exploration of the transformative impact of LLMs on IR across various dimensions. We have organized existing approaches into distinct categories based on their functions: query rewriting, retrieval, reranking, and reader modules. In the domain of query rewriting, LLMs have demonstrated their effectiveness in understanding ambiguous or multi-faceted queries, enhancing the accuracy of intent identification. In the context of retrieval, LLMs have improved retrieval accuracy by enabling more nuanced matching between queries and documents, considering context as well. Within the reranking realm, LLM-enhanced models consider more fine-grained linguistic nuances when re-ordering results. The incorporation of reader modules in IR systems represents a significant step towards generating comprehensive responses instead of mere document lists. The integration of LLMs into IR systems has brought about a fundamental change in how users engage with information and knowledge. From query rewriting to retrieval, reranking, and reader modules, LLMs have enriched each aspect of the IR process with advanced linguistic comprehension, semantic representation, and context-sensitive handling. As this field continues to progress, the journey of LLMs in IR portends a future characterized by more personalized, precise, and user-centric search encounters.
This survey focuses on reviewing recent studies of applying LLMs to different information retrieval components. Beyond this, a more significant problem brought by the appearance of LLMs is: is the conventional IR framework necessary in the era of LLMs? For example, traditional IR aims to return a ranking list of documents that are relevant to issued queries. However, the development of generative language models has introduced a novel paradigm: the direct generation of answers to input questions. Furthermore, according to a recent perspective paper [53], IR might evolve into a fundamental service for diverse systems. For example, in a multi-agent simulation system [162], an IR component can be used for memory recall. This implies that there will be many new challenges in future IR.